You drive to the mailbox down the road at 35 mph, slip a letter in, turn around, and drive back at 25 mph. What is your average speed? (Note that it is not 30 mph, and you do not need to know the distance.) I tried this but don't know where I am going wrong.
T (going)= D/35
and for the return trip
T(returning) = D/25
The total time for the trip is T(going) +T(returning) = D/35 + D/25 and you can find the average speed by dividing the total distance 2*D, by the total time:
V(average) = 2D/(D/35+D/25)
V(average)= 1/(1/35 + 1/25)

V(average)= (1*60/1)/(1/60*60/1)
V(average)= 60 mph I know this answer is wrong. It doesn't make sense that it should be the combined speeds.

Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

You drive to the mailbox down the road at 35 mph, slip a letter in, turn around, and drive back at 25 mph. What is your average speed? (Note that it is not 30 mph, and you do not need to know the distance.) I tried this but don't know where I am going wrong.
T (going)= D/35
and for the return trip
T(returning) = D/25
The total time for the trip is T(going) +T(returning) = D/35 + D/25 and you can find the average speed by dividing the total distance 2*D, by the total time:
V(average) = 2D/(D/35+D/25)
V(average)= 1/(1/35 + 1/25) **** You changed the NUM to 1, should be 2 *****
Avg = 2/(1/35 + 1/25)
Avg = 2/(5/175 + 7/175) = 2/(12/175)
Avg = 175/6
Avg = ~ 29.1667 mph
-------------------
This is the basis of the Michelson-Morley experiment that disproved the existence of "the ether" (not exactly the same, but similar).
V(average)= (1*60/1)/(1/60*60/1)
V(average)= 60 mph I know this answer is wrong. It doesn't make sense that it should be the combined speeds.

Answer by solver91311(24713)


Actually you were on the right track until you wrote:



Which is incorrect. It should have been:



Since you were only taking out a factor of

And then how you got from there to:



Is completely beyond me. So lets go back to:



The LCD is 175, so:





John